OpenStudy (anonymous):

The position of a particle in meters is given by, x(t)=16t-3.0t^3 where the time t is in seconds. At what time is the particle momentarily at rest ?

OpenStudy (anonymous):

x(t) must be equal to 0.differentiating 16t-3t^3 and putting it equal to zero we get 16-9t^2=0 9t^2=16 t^2=16/9 t=4/3

OpenStudy (anonymous):

The height of a helicopter above the ground is given by h = 3.00t^3 where h is in meters and t is in seconds . After 2.00s, the helicopter releases a small mailbag . How long after its release does the mailbag reaches the ground ??